OpenStudy (anonymous):

The angle that a waterskiing ramp forms with the surface of the water is 1.5, and the ramp rises 5 ft. Approximately how long is the ramp?

OpenStudy (anonymous):

\[R \sin(\theta)=y\] \[R=\frac{ y }{ \sin(\theta) }\] where y is the height and \[\theta\] is your angle